The sum of five consecutive even numbers is equal to 170.What is the sum of the second largest number amongst them and the square of the smallest number amongst them together ?
Answer by suruman(21)

Let the five consecutive even numbers be :
x-4,x-2,x,x+2,x+4
Sum of the five consecutive even numbers = 170.
There fore,
{ (x-4)+(x-2)+(x)+(x+2)+(x+4) } = 170.
5*x = 170
x = 170/5 = 34
Hence the consecutive even numbers are : 30,32,34,36 and 38.
Smallest number among them is 30.
Second largest number among them is 32.
Hence, sum of second largest number and the square of the smallest number is :
30*30 + 32 = 932.